The empheq package^{∗} Emphasizing equations in L^{A}TEX 2ε

Morten Høgholm

2004/10/10

Abstract The empheq package can best be described as a visual markup extension

to amsmath. In short it offers a) a multi line equivalent of \boxed from amsmath and b) a way to produce arbitrary delimiters that span entire math displays.

Contents

I Basic user’s guide 3

1 Basic use of the package 3 1.1 Using the empheq environment . . . . . . . . . . . . . . . . . . . . 3 1.2 Markup instructions . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 What won’t work in the empheq environment . . . . . . . . . . . . 6 1.4 Special delimiters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2 Taking things a little further 7

2.1 Loading the package . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1.1 Compatibility with the old version . . . . . . . . . . . . . . 8 2.2 Package options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.2.1 Using multline . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.2 The overload option . . . . . . . . . . . . . . . . . . . . . 10

2.3 A note on boxed displays . . . . . . . . . . . . . . . . . . . . . . . 12

3 Support for other packages 13 3.1 Support for fancybox . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.2 Support for ntheorem . . . . . . . . . . . . . . . . . . . . . . . . . . 15

II Advanced user’s guide 17

4 Delimiters revisited 17 4.1 Creating your own delimiters . . . . . . . . . . . . . . . . . . . . . 17 4.2 Fine-tuning of delimiters . . . . . . . . . . . . . . . . . . . . . . . . 17 4.3 Scaling material yourself . . . . . . . . . . . . . . . . . . . . . . . . 18 5 Creating something new 19

∗This file has version number v2.11a, last revised 2004/10/10.

5.1 New empheq-like environments . . . . . . . . . . . . . . . . . . . . 19 5.2 Creating fancy boxes . . . . . . . . . . . . . . . . . . . . . . . . . . 20

6 Contact information 21

Part I

Basic user’s guide

There can be little doubt that the de facto standard for mathematical typesetting in L^{A}TEX is the amsmath package. For the creation of empheq, a visual markup package for use in math, it made perfect sense to have amsmath as the backbone.

The main idea of empheq is to maintain the familiar syntax of the amsmath environments while still providing an easy way of specifying markup instructions. This manual is plastered with examples showing just how.

1 Basic use of the package

So what is it empheq does? Well, it allows you to produce displays like this:

X = Y ⇒ \begin{cases}A_{1} & = b_{1} & c_{1} & = d_{1} & e_{1} & = f_{1} \\ A_{2} & = b_{2} & c_{2} & = d_{2} & e_{2} & = f_{2} \\ A_{3} & = b_{3} & c_{3} & = d_{3} & e_{3} & = f_{3} \end{cases}!  Asillytag (1a) (1b)

In short empheq enables the user to put things on every side of the display without said user having to worry about what happens to the equation numbers. For example you can now have a display containing multiple lines and still get the effect of the \boxed command from amsmath.

1.1 Using the empheq environment

The package defines a single environment empheq and the usage is kind of straight forward:

\begin{empheq}[⟨markup instructions⟩]{⟨AMS env˙ name⟩}

⟨contents of AMS environment⟩

 \end{empheq}

A first minimal example file would then be something like

    \documentclass{minimal}
    \usepackage{empheq}
    \begin{document}
    \begin{empheq}{align*}
      a&=b \tag{*}\\
      E&=mc^2 + \int_a^a x\, dx
    \end{empheq}
    \end{document}

This gives the following display:

\begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned}  (*)

Environment Usage  Environment Usage

     equation  {equation}  equation*  {equation*}
     align  {align}  align*  {align*}
     gather  {gather}  gather*  {gather*}
     flalign  {flalign}  flalign*  {flalign*}
     alignat  {alignat=⟨cols⟩} alignat*  {alignat*=⟨cols⟩}
     multline  {multline}  multline*  {multline*}

Table 1: The supported amsmath environments

Maybe not the most impressive example, but as you can see the contents of the environment is exactly the same as for the regular align* evironment from amsmath. The rest of the AMS environments are chosen the same way by typing the name as the mandatory argument of empheq with the exception of the alignat environment. For this you have to specify the number of columns as shown below.

  \begin{empheq}{alignat=2}
    a &= b &\quad c &= d \\
    \text{this} &= \text{that} &\quad \mathit{fish}&\neq fish
  \end{empheq}

a = b  c = d
this = that fish ≠ fish

{(2) \atop (3)}

To choose the starred version of alignat, simply type {alignat*=2} instead in the above example.

The supported AMS environments are listed in Table 1. Not supported is the standard L^{A}TEX eqnarray environment as it is fundamentally deficient.^{1}

1.2 Markup instructions

The optional argument of the empheq environment will take markup instructions consisting of a ⟨key⟩ = ⟨value⟩ list of assignments. There are currently five such keys (a sixth is added for naming consistency). They’re shown in Table 2 on the next page.

left

The key left is for material put on the left side of the display. The material is typeset in math mode and centered vertically.

  \begin{empheq}[left=L\Rightarrow]{align}
    a&=b\\
    E&=mc^2 + \int_a^a x\, dx
  \end{empheq}

L ⇒ \begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned}  {(4) \atop (5)}

1See for instance http://www.tex.ac.uk/cgi-bin/texfaq2html?label=eqnarray

Key  Usage  Additional Info
box  box=⟨box command⟩
innerbox innerbox=⟨box command⟩

left left=⟨math material⟩ Use \text{...} if you need

text material.

right right=⟨math material⟩ Use \text{...} if you need

text material.

outerbox outerbox=⟨box command⟩ Alias for box.

marginbox marginbox=⟨box command⟩ Contents must be horizontally centered; can only be used in

                                       [fleqn] mode.

Table 2: The six keys for the optional argument of the empheq environment

right

As there is a left key it hopefully comes as no surprise that there is a companion right key for typesetting material on the right side of the display.

  \begin{empheq}[right=\Leftarrow R]{align}
    a&=b\\
    E&=mc^2 + \int_a^a x\, dx
  \end{empheq}

\begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned} ⇐ R  {(6) \atop (7)}

box outerbox

The key box specifies the kind of box you would like to put around the display. It can be any kind of box, as long as the contents of the box is situated on the baseline like in a \fbox .

  \begin{empheq}[box=\fbox]{align}
    a&=b\\
    E&=mc^2 + \int_a^a x\, dx
  \end{empheq}

\begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned}  {(8) \atop (9)}

The key outerbox is an alias for box and is added for naming consistency with the key described below.

innerbox

There is also an innerbox key. It is not very interesting unless you use one of the other keys.

  \begin{empheq}[innerbox=\fbox,
                 left=L\Rightarrow]{align}
    a&=b\\
    E&=mc^2 + \int_a^a x\, dx
  \end{empheq}

L ⇒ \begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned}  {(10) \atop (11)}

marginbox

The last key is the marginbox key. If you typeset your math in [fleqn] mode^{2} you

Feature request by Uwe Siart 2003/12/08

may want the math display and not the outer box to align at the left margin (or rather: the indentation). If you make sure the contents of the outer box is centered horizontally inside the box, marginbox will align it properly. You shouldn’t set both box and marginbox at the same time, as this is surely not what you want and the package will silently use the last one in the list.

Warning: keyval treats commas and equal signs as separator and assignment signs which in turn means that if you want to typeset either of them, they must—and believe me, it’s really important—be enclosed in braces. In short:

     Good [left={(A,B)=(1,0)}]
     Bad [left=(A,B)=(1,0)]

\empheqset{⟨markup instructions⟩}

You can also set the keys globally^{3} with the command \empheqset. This means

Feature request by Uwe Siart 2004/01/13

that

    \empheqset{marginbox=\psframebox}

will force a box=\psframebox in all occurrences of the empheq environment, but an explicitly given box or marginbox will override this setting. You can only use the keys listed in this section as arguments to \empheqset; not the AMS environments.

1.3 What won’t work in the empheq environment

 \intertext{⟨text⟩}
 \displaybreak[⟨num⟩]

Now we’ve seen some of the things that work in the empheq environment but we also have to take note of what won’t work. As this package provides a way to box multi line math displays it can come as no surprise that using either \intertext or \displaybreak inside the empheq environment makes no sense. Should you however happen to try them anyway, you’ll experience that \intertext issues an error message and \displaybreak issues a warning. The reason only \intertext gives an error message is that you’ll get output very different from what you expect, and that is not the case (so much at least) with \displaybreak. But of course, you’ll never see either message because you read the manual!

2Sorry, but I can’t show you because this manual features centered math displays.

3Or rather: outside the scope of the empheq environment. The settings will still obey scoping rules.

1.4 Special delimiters

As you’ve seen a few pages back, it’s possible to add material on both sides of the math display. When doing so you often need a delimiter that scales to fit the entire display, so it comes as no surprise that empheq provides such delimiters.

\empheql⟨delim˙ name⟩ \empheqbigl⟨delim˙ name⟩ \empheqr⟨delim˙ name⟩ \empheqbigr⟨delim˙ name⟩

For example one might need a large brace like the one in the cases environment:

  \begin{empheq}[left=\empheqlbrace, right=\empheqrbrace]{align}
    E&=mc^2 \\
    Y&= \sum_{n=1}^\infty \frac{1}{n^2}
  \end{empheq}

\begin{cases}E & = mc^{2} \\ Y & = ∑_{n=1}^{∞} \frac{1}{n^{2}} \end{cases} \}  {(12) \atop (13)}

The naming scheme is \empheq⟨delimiter name⟩ and \empheqbig⟨delimiter name⟩. Thus \empheqrbrace produces an auto-scaling right brace with same size as the math display while \empheqbigrbrace produces an even bigger version that spans the inner box plus the math display inside it. For a complete list of supported delimiters see Table 3 on the following page.

2 Taking things a little further

So far we have covered the basic functions of empheq: the markup instructions and the predefined auto-scaling delimiters. There is more to empheq however, and we’ll cover that in this section.

2.1 Loading the package

The package has two main requirements: One is the mathtools^{4} package which provides empheq with a lot of necessary tools for doing its thing. The other one is amsmath of course, and to make the loading procedure as easy as possible, you can simply substitute

  \usepackage[leqno,fleqn,intlimits]{amsmath}

4By the same author and is distributed with empheq. See its documentation for more information.

Delimiter

Original  Normal  Bigger  Symbol

\lbrace  \empheqlbrace  \empheqbiglbrace  {
\rbrace  \empheqrbrace  \empheqbigrbrace  }
\lbrack  \empheqlbrack  \empheqbiglbrack  [
\rbrack  \empheqrbrack  \empheqbigrbrack  ]
\langle  \empheqlangle  \empheqbiglangle  ⟨
\rangle  \empheqrangle  \empheqbigrangle  ⟩
\lparen  \empheqlparen  \empheqbiglparen  (
\rparen  \empheqrparen  \empheqbigrparen  )
\lvert   \empheqlvert   \empheqbiglvert  |
\rvert   \empheqrvert   \empheqbigrvert  |
\lVert   \empheqlVert   \empheqbiglVert  ∥
\rVert   \empheqrVert   \empheqbigrVert  ∥
\lfloor  \empheqlfloor  \empheqbiglfloor  ⌊
\rfloor  \empheqrfloor  \empheqbigrfloor  ⌋
\lceil   \empheqlceil   \empheqbiglceil  ⌈
\rceil   \empheqrceil   \empheqbigrceil  ⌉

Table 3: The supported auto-scaling delimiters in empheq

with

\usepackage[leqno,fleqn,intlimits]{empheq}

empheq makes sure that the amsmath options are passed on and loaded by amsmath. The same goes for the options provided by mathtools.

2.1.1 Compatibility with the old version

The current version (v2.11a) of empheq is incompatible with versions prior to and including 0.7d. If you have documents produced with versions prior to 0.7d, then you have to load the package empheq07 instead. empheq07 now exists as a separate package package with its own documentation, but I strongly recommend switching to the new version because it is so much better.

2.2 Package options

In addition to the amsmath options, empheq itself provides a string of options listed in Table 4 on the next page.

ntheorem

The ntheorem package is supported by means of the ntheorem option. In order for this to work properly, an amsmath bug^{5} is fixed. For more information on this option see § 3.2 on page 15.

5See the L^{A}TEX Bugs Database http://www.latex-project.org/cgi-bin/ltxbugs2html under AMSL^{A}TEX problem report 3624.

Option  Short description

overload Lets you use the empheq visual markup exten-

sions transparently in an existing document.

overload2  A wild version of overload. Use with care.

ntheorem A support module for users of ntheorem and its

                    thmmarks and thref options.

newmultline With v2.10 of empheq the syntax for multline

and multlined has been improved immensely.

oldmultline Support for the somewhat strange syntax for

                    multline and multlined in v2.00.

Table 4: Supported options in the empheq package

2.2.1 Using multline

 newmultline oldmultline

Those familiar with v2.00 of this package will recognize the somewhat weird syntax for using multline and multlined. The good news is that since v2.10 there is now a much improved syntax available. The option newmultline (default) selects this better interface, while the old interface is provided for compatibility reasons by the oldmultline option.

 \shoveleft[⟨dimen⟩]{⟨arg⟩}
 \shoveright[⟨dimen⟩]{⟨arg⟩}

With the new interface you also get an extended syntax for \shoveleft and \shoveright as shown in the example below.

  \begin{empheq}{multline}
    \framebox[.65\columnwidth]{First line} \\
    \framebox[.5\columnwidth]{Second line} \\
    \shoveleft{L+E+F+T}         \\
    \shoveright{R+I+G+H+T}         \\
    \shoveleft[1cm]{L+E+F+T}         \\
    \shoveright[\widthof{$R+I+G+H+T$}]{R+I+G+H+T}         \\
    \framebox[.65\columnwidth]{Last line}
  \end{empheq}

First line

Second line

L + E + F + T

R + I + G + H + T
L + E + F + T
R + I + G + H + T

Last line (14)

There are however a few differences in the output between the original amsmath version of multline and the one empheq provides. In amsmath a centered line in multline is centered on the page without taking into account the \multlinegap, \multlinetaggap, and the tag width. Thus amsmath can sometimes give you horrible output without giving you any warning as shown below.

  \begin{multline}
    \framebox[.65\columnwidth]{First line} \\
    \framebox[.9\columnwidth]{Loooong line} \\
    \framebox[.65\columnwidth]{Last line} \tag{wide tag}
  \end{multline}

First line

Loooong line

Last line  (wide tag)

In empheq these parameters are taken into account, so the same input will produce this in empheq:

  \begin{empheq}{multline}
    \framebox[.65\columnwidth]{First line} \\
    \framebox[.9\columnwidth]{Loooong line} \\
    \framebox[.65\columnwidth]{Last line} \tag{wide tag}
  \end{empheq}

First line Loooong line Last line  (wide tag)

This results in an

    Overfull \hbox (30.03783pt too wide) in paragraph ...

message in the log file, indicating a visual problem. I think this behavior is more sensible than the one the original multline environment provides.

 \begin{MTmultlined}[⟨pos⟩][⟨width⟩] ⟨contents⟩ \end{MTmultlined}

When you choose the oldmultline option you may still want to use the much improved multlined environment defined in mathtools but unfortunately there will be a name clash. Instead you can access it under the name MTmultlined.

2.2.2 The overload option

 overload overload2

The overload option redefines the original AMS environments so that they take

Feature request by Lars Madsen 2004/03/25

an optional argument.

\begin{⟨AMS env˙ name⟩}[⟨markup instructions⟩] ⟨contents of AMS environment⟩ \end{⟨AMS env˙ name⟩}

For example

    \begin{gather}[box=\widefbox]
      a = b
    \end{gather}

is then actually short for

    \begin{empheq}[box=\widefbox]{gather}
      a = b
    \end{empheq}

All the AMS environments are supported by this option except for the pseudo environment \[ ... \] (it’s not really an environment), ^{6} because a) you don’t really need markup instructions for a one line display, and b) would you really like a syntax like

    \[[box=\fbox] a=b \]

where its difficult to see whether or not there is a typo? Choosing overload is meant for people who don’t want to change their entire document into empheq syntax. I have no problems with that; just be careful when you fiddle with \empheqset as it will affect all math displays!

There is of course a catch (well, two actually): These redefined environments don’t run as fast as the regular ones (about three times as slow), but in this day and age I seriously doubt you’ll be able to tell the difference unless you have a vast number of equations. The other catch is that you cannot use \intertext and \displaybreak as described earlier. If you find yourself wanting to use one of these features in say, an align environment, you have to use the original align environment. Luckily it is available if you call it like this:

\begin{AmS⟨AMS env˙ name⟩}

⟨contents of AMS environment⟩ \end{AmS⟨AMS env˙ name⟩}

For example the original align environment could be selected with

    \begin{AmSalign}
      ...
    \end{AmSalign}

These original versions with prefix AmS exist for all the environments.

The option overload2 activates the overloading feature for the pseudo environment \[ ... \], although I doubt you’ll find it useful. Beware that this definition is fragile unless you have ε-TEX as L^{A}TEX engine.^{7} Not surprisingly

overload2 activates overload.

Before you get all excited about these options, you should take note of some aspects of centered math displays in amsmath. In certain circumstances truly centering the display is not always the best solution as in this example:

6With the option oldmultline the multline and multline* environments aren’t supported either because their syntax in empheq then differ from their syntax in amsmath.

7ε-TEX has been the default engine for L^{A}TEX in most major distributions since 2003.

Wide math display; not adjusted Wide Tag

Instead centering the display in the available space works pretty well:

Wide math display; adjusted

Wide Tag

This is what amsmath does normally, but in the gather environment it does it on a per line basis. This means that we get results like

Wide math display; adjusted

Wide Tag

Wide math display; not adjusted

(i)

whether we like it or not. Inside the empheq environment we need to have the same adjustment for all lines else the boxing process will not work properly, so when activating the overload option the above two-line gather will instead look like this:

Wide math display; adjusted

Wide math display; adjusted

Wide Tag i

I leave it to you to choose whether or not this is better (I think it is better).

2.3 A note on boxed displays

When browsing a 400+ pages textbook with at least twice the number of displayed formulae, some of them are surely more important than others. Thus the author of the book (in cooperation with the designer) should make sure that such formulae are easily found again i.e., they should be easily distinguishable from the rest of the pack. One way of doing this is putting the formula into a box which is something we’ve seen empheq being capable of. There are however a few things to keep in mind:

• Don’t overdo it. If you do it on half of them there’s no point in doing it at all. I don’t see much reason for applying this technique to more than 10% of the formulae.

• Choose the type of box carefully. You want to draw attention to it so it might as well look good.

The latter point can be illustrated by defining a macro similar to \fbox only with a little more space to the right and left of the argument.

  \newcommand*\widefbox[1]{\fbox{\hspace{1em}#1\hspace{1em}}}

If we replace the \fbox from the box example before, we get this display:

  \begin{empheq}[box=\widefbox]{align}
    a&=b\\
    E&=mc^2 + \int_a^a x\, dx
  \end{empheq}

\begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned}  {(15) \atop (16)}

Compare it with

\begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned}  {(17) \atop (18)}

and see which one you prefer.

Similarly one might be tempted to use colored boxes:

  \definecolor{myblue}{rgb}{.8, .8, 1}
  \newcommand*\mybluebox[1]{%
    \colorbox{myblue}{\hspace{1em}#1\hspace{1em}}}

We know the drill by now.

  \begin{empheq}[box=\mybluebox]{align}
    a&=b\\
    E&=mc^2 + \int_a^a x\, dx
  \end{empheq}

\begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned}  {(19) \atop (20)}

There is more on boxes later in this manual. If you’re into the fancybox package then remember to read § 3.1 on the next page.

3 Support for other packages

With the multitude of packages for L^{A}TEX, it is not always easy to be a PWWO package (“Plays Well With Others”), but empheq tries really hard to do so. This section lists the packages where empheq has to provide workarounds and they can be divided into two categories.

Compatibility Some packages affect L^{A}TEX’ labelling mechanism and since empheq internally has to turn off labels and related code temporarily, hooks must be provided for these packages. Examples of such packages are hyperref and showkeys.

Enhancements Other packages provide useful features that for some reason may not work directly or optimally with empheq. In these cases the problematic commands are redefined so that they not only work with empheq, but also give the same if not better output. An example of this would be the \shadowbox command from fancybox (described below).

3.1 Support for fancybox

The fancybox package provides various boxes and you can use them with empheq as well. Here’s \ovalbox:

  \begin{empheq}[box=\ovalbox]{align}
      E&=mc^2 \\
      Y&= \sum_{n=1}^\infty \frac{1}{n^2}
  \end{empheq}

[figure: \begin{aligned}E & = mc^{2} \\ Y & = ∑_{n=1}^{∞} \frac{1}{n^{2}} \end{aligned}]  {(21) \atop (22)}

 \shadowbox*{⟨arg⟩}
 shadowcolor

The only problem with using the fancybox boxes in conjunction with the empheq environment is \shadowbox, as this macro typesets the shadow on the baseline and for the equation numbers to be placed correctly, the box command must set

its argument on the baseline. Thus the normal \shadowbox produces ⟨arg⟩ , but we need this instead: ⟨arg⟩ . Therefore empheq will detect if fancybox is loaded

and in that case it’ll enhance \shadowbox in two ways:

• It defines a starred version \shadowbox* which typesets its argument on the baseline.

• The color shadowcolor is introduced. The default color is black.

  \definecolor{shadowcolor}{rgb}{0,.5,.5}
  \setlength\shadowsize{2pt}
  Line of text for \shadowbox*{testing of \Verb|f#&%\| verbatim} and
  showing the shadow color.

Line of text for testing of f#&%\ verbatim and showing the shadow color.

The point is that if you want a \shadowbox around your math display then you must use the starred version:

  \begin{empheq}[box=\shadowbox*]{align}
      E&=mc^2 \\
      Y&= \sum_{n=1}^\infty \frac{1}{n^2}
  \end{empheq}

\begin{aligned}E & = mc^{2} \\ Y & = ∑_{n=1}^{∞} \frac{1}{n^{2}} \end{aligned}  {(23) \atop (24)}

See § 5.2 on page 20 if you want to make a fancy box yourself.

Math Height Math Depth Tag height Tag depth  Marker

Figure 1: Marker placement in [leqno] mode

[figure: Math Height; Math Depth; Tag height; Tag depth; Marker; > δ; Math Height; Math Depth; Tag height; Tag depth; Marker; = δ]

Figure 2: The two possible situations for end marks in [reqno] mode

3.2 Support for ntheorem

 ntheorem

The ntheorem package is supported by the means of the ntheorem option of empheq. This loads a set of extra macros which fixes various compatibility problems between ntheorem and amsmath and furthermore introduces special (internal) macros for optimum positioning of end-of-theorem markers, while retaining a user friendly interface.

When you want to use the automatic end-of-theorem marker mechanism from ntheorem you sometimes run into problems as you would want the marker to be placed aligned at the bottom of the math display but still keeping the tags in their proper place. In [leqno] mode this is not that much of a problem as the tags and the marker are set on either side of the math display like in Figure 1.

 \mintagvsep

Unfortunately things are not this easy in [reqno] mode. There are two possible situations as shown in Figure 2. If possible we want the bottom of the marker to be aligned with the bottom of the math display, but at the same time we want to ensure a minimum vertical separation δ between marker and tag. In empheq this is controlled by the length parameter \mintagvsep which by default is 5 pt.

The good news is that this is where empheq sets in. Basically all you need to do is to use the empheq environment to type set your equations inside the theorem environments:

    \begin{Theorem}

    Some text at first and then a math display
    \begin{empheq}{align}
      a&=b\\
      E&=mc^2 + \int_a^a x\, dx
    \end{empheq}
    \end{Theorem}

Then the tag placement and the end-of-theorem marker will be set properly (after a couple of runs as usual).

To ensure the correct outcome remember to load the packages like this:

    \usepackage[ntheorem]{empheq} % this loads amsmath as well
    \usepackage[thmmarks,amsmath]{ntheorem}

Remember that if you use the overload option you can just use the regular math environments to get the desired result with the end mark. The exception is the \[ ... \] environment which will only work if you use the option overload2.

Part II

Advanced user’s guide

The empheq package has more to offer than you have seen, but I found some of the functionality so dangerous it was best to hide it for a little while. The commands you have encountered so far in this manual all have one thing in common: They have only lower case letters in their names. Now it’s time to reveal those that have mixed-case names, thus implying that you should take great care when using them.

4 Delimiters revisited

Let’s go back to delimiters, shall we?

4.1 Creating your own delimiters

As a convenience for the user, this interface is extended in a more general way so that it is also possible to declare delimiters with the following two commands:

\DeclareLeftDelimiter[⟨space adjustment⟩]{⟨delimiter⟩} \DeclareRightDelimiter[⟨space adjustment⟩]{⟨delimiter⟩}

While empheq provides auto-scaling versions of the most common delimiters, you may sometimes want some new ones. Say for instance you have loaded the stmaryrd package and you want to use the double bracket commands \llbracket and \rrbracket with empheq. Then you simply do this:

    \DeclareLeftDelimiter{\llbracket}
    \DeclareRightDelimiter{\rrbracket}
This defines the new delimiters \empheqllbracket, \empheqbigllbracket, \empheqrrbracket
and \empheqbigrrbracket.

You can use \big... delimiters as well if you don’t like the automatic scaling. There are however ways to fine-tune if you really want it.

4.2 Fine-tuning of delimiters

TEX provides two primitives to control the scaling of delimiters produced with \left and \right, namely the dimension \delimitershortfall (denoted by δ here) and the integer \delimiterfactor (f). The idea is that the sub-formula inside the \left-\right pair is to be vertically centered and given its height h_{1} and its depth h_{2} we want to produce a delimiter with total height h, where h = 2 max(h_{1}, h_{2}). TEX’ rules on this is that the minimum delimiter size h_{min} must meet the requirements

h_{min} ≥ h\frac{f}{1000} ∧ h_{min} ≥ h − δ

L^{A}TEX sets \delimitershortfall = 5.0pt and \delimiterfactor = 901, but in

our case we will almost always want a delimiter that spans the entire \left-\right pair (the math display), thus a change of these settings is needed.

 \EmphEqdelimitershortfall
 \EmphEqdelimiterfactor

However it is a bad idea to just change these two settings without thinking of the effect it’ll have on the rest of the mathematics in your document. Therefore empheq provides the parameters \EmphEqdelimitershortfall (default setting is 1.0pt) and \EmphEqdelimiterfactor (default is 950) to cater for this.

4.3 Scaling material yourself

The attentive reader may have noticed that I still haven’t revealed how I managed to get the right size of that big exclamation mark in the first example of this manual, so I guess it’s about time.

 \EmphEqdisplayheight
 \EmphEqdisplaydepth

The height of the math display plus the surrounding inner box is given by

\EmphEqdisplayheight and the depth by \EmphEqdisplaydepth. These can be

used when constructing something that should scale to fit the display and can’t be done with “auto-scaling” commands like \vrule or \left or \right. There is a catch however: If you do any horizontal resizing of the material you want to scale, then you must specify the width manually:^{8}

  \begin{empheq}[
    box=\colorbox{myblue},
    right={\;
           \makebox[.9em]{%
            $\raisebox{-.5\totalheight+\fontdimen22\textfont2}
              {\resizebox{!}{%
                \EmphEqdisplayheight+\EmphEqdisplaydepth}{!}}$%
           }},
    left={X=Y\Rightarrow}
     ]{alignat=3}
    A_1&=b_1 & \qquad c_1&=d_1 & \qquad e_1&= f_1
         \tag*{\_A\raisebox{1ex}{silly}\raisebox{-1ex}{tag}\_}\\
    A_2&=b_2 & \qquad c_2&=d_2 & \qquad  e_2&= f_2 \\
    A_3&=b_3 & \qquad c_3&=d_3 & \qquad  e_3&= f_3
  \end{empheq}

Here’s another example. We want to be able to put a \parbox with some descriptive text top-aligned on the side of the display.

  \begin{empheq}[
    left={\parbox[c][\EmphEqdisplayheight+\EmphEqdisplaydepth][t]
           {4.5cm}
           {You may find this kind of description useful.}\enspace}]
      {align}
    a&=\int_0^1 x\,dx +\frac{foo + bar}{baz}\\
    E&= mc^2

8The \fontdimen22\textfont2 in the ⟨length⟩ argument of \raisebox is the math axis of the font. It is needed in order to get the vertical positioning of the oversized exclamation mark just right.

  \end{empheq}

You may find this kind of de- scription useful. \begin{aligned}a & = ∫_{0}^{1} x dx + \frac{foo + bar}{baz} \\ E & = mc^{2} \end{aligned}  {(25) \atop (26)}

5 Creating something new

You can create your own environments and boxes to go with empheq, but there are certain things that must be fulfilled to get it to work properly.

5.1 New empheq-like environments

 EmphEqMainEnv
 \EmphEqMainEnv
 \endEmphEqMainEnv

Although the real work in empheq is done by an environment (EmphEqMainEnv), you cannot use it like that yourself when you want to define an empheq-like environment. Due to technical reasons (see [4]), you have to use a slightly different syntax:

\newenvironment {⟨env˙ name⟩} [⟨num⟩] [⟨default⟩]

   {⟨other commands ⟩ \EmpheqMainEnv}
   {\endEmpheqMainEnv}

In the above environment the ⟨other commands⟩ part must contain certain things before it’ll work.

EmphEqEnv EmphEqOpt

Before your homemade environment will work you must set at least one key. The key family EmphEqOpt is for all the markup instructions like left and box, while EmphEqEnv controls the type of AMS environment. You must set the EmphEqEnv family before \EmhpEqMainEnv else you’ll get an error message.

One example could be something like this:

  \newenvironment{important}[2][]{%
    \setkeys{EmphEqEnv}{#2}%
    \setkeys{EmphEqOpt}{box=\mybluebox,#1}%
    \EmphEqMainEnv}%
    {\endEmphEqMainEnv}

Thus

  \begin{important}{gather}
    a = b + c +d \\ e = f
  \end{important}

produces

\begin{matrix}a = b + c + d \\ e = f \end{matrix}  {(27) \atop (28)}

while

  \begin{important}[left={A=B \Rightarrow \empheqlbrace}]{alignat=2}
    a &= b &\quad c &= d \\
    \text{this} &= \text{that} &\quad \mathit{fish}&\neq fish
  \end{important}

produces

A = B ⇒ \begin{cases}a & = b & c & = d \\ this & = that & fish & ≠ fish \end{cases}  {(29) \atop (30)}

5.2 Creating fancy boxes

As a final example I will show you how to create complicated displays involving (vertically) asymmetrical boxes like \shadowbox*. In order to get the correct output, the contents of the box must be placed on the baseline as in \fbox . In this example we want to put a set of equations into a bright yellow box and then add another box with some explanatory text at the top of the yellow box making them overlap.

First we define the colors used and allocate the boxes. We could probably use scratch boxes, but this is safer.

  \definecolor{shadecolor}{cmyk}{0,0,0.41,0}
  \definecolor{light-blue}{cmyk}{0.25,0,0,0}
  \newsavebox{\mysaveboxM} % M for math
  \newsavebox{\mysaveboxT} % T for text

Save the display body in \mysaveboxM and the text argument in \mysaveboxT.

  \newcommand*\Garybox[2][Example]{%
    \sbox{\mysaveboxM}{#2}%
    \sbox{\mysaveboxT}{\fcolorbox{black}{light-blue}{#1}}%

Then typeset the math display in a \parbox where we control the height and save

it in \mysaveboxM.
    \sbox{\mysaveboxM}{%
      \parbox[b][\ht\mysaveboxM+.5\ht\mysaveboxT+.5\dp\mysaveboxT][b]{%
        \wd\mysaveboxM}{#2}%
    }%

We put it into the colored box with the desired width.

    \sbox{\mysaveboxM}{%
      \fcolorbox{black}{shadecolor}{%
        \makebox[\linewidth-10em]{\usebox{\mysaveboxM}}%
      }%
    }%

Then finally we get to the real typesetting. We just insert the math display and then make a box of zero width right next to it. In that box we lift the text argument and center it at the top of the display.

    \usebox{\mysaveboxM}%
    \makebox[0pt][r]{%
      \makebox[\wd\mysaveboxM][c]{%
        \raisebox{\ht\mysaveboxM-0.5\ht\mysaveboxT
                  +0.5\dp\mysaveboxT-0.5\fboxrule}{\usebox{\mysaveboxT}}%
      }%
    }%
  }

Let’s see what it looks like.

  \begin{empheq}[box=\Garybox]{align}
    \sum \mathbf{F} &= \mathbf{0} \\
    \sum F_{x} \,\mathbf{i} + \sum F_{y} \,\mathbf{j}
      + \sum F_{z}  \,\mathbf{k} &= \mathbf{0} \\
    \sum F_{x} &= 0 \\ \sum F_{y} &= 0 \\ \sum F_{z} &= 0
  \end{empheq}

\begin{aligned}∑ F & = 0 \\ ∑ F_{x} i + ∑ F_{y} j + ∑ F_{z} k & = 0 \\ ∑ F_{x} & = 0 \\ ∑ F_{y} & = 0 \\ ∑ F_{z} & = 0 \end{aligned} Example  (31a) (31b) (31c) (31d) (31e)

If we use the optional argument of \Garybox we have to enclose the entire argument of box in braces because we’re in empheq’s optional argument (as described in [1, page 167]):

  \begin{empheq}[box={\Garybox[Same old example again]}]{align}
    a&=b\\
    E&=mc^2 + \int_a^a x\, dx
  \end{empheq}

\begin{aligned}a & = b \\ E & = mc^{2} + ∫_{a}^{a} x dx \end{aligned} Same old example again  {(31f) \atop (31g)}

6 Contact information

Should you have any feature request, suggestions, or bug reports then feel free to contact me at

     moho01ab@student.cbs.dk

Please put ‘empheq’ in the subject line.

Contributors

• In November 2002 Lars Madsen (daleif@imf.au.dk) asked for some features that wasn’t readily available with amsmath or any other package. This was the start of empheq.

• Gary Gray (gray@engr.psu.edu) gave me some bug reports on the old version (v0.7) which convinced me to rewrite the package completely. The \Garybox example is inspired by a wish from Gary.

• Uwe Siart (uwe.siart@tum.de) has been a thorough beta-tester on v1.00 of the package.

• Bernard Alfonsi (alfonsi@math.u-psud.fr) reported a problem with the ntheorem option. It turned out to be a bug in the ntheorem package itself but it’s fixed in the option.

• I have received reports on weird behavior from empheq when used with the color package on some dvi-viewers, which is caused by the lack of color support in those dvi-viewers. Both Andrew B. Collier (colliera@ukzn.ac.za) and André M. de Roos (aroos@science.uva.nl) have notified me of this. The solution is to either upgrade you dvi-viewer if possible or convert the document in question to PDF or PS.

Thank you all.

References

[1] Leslie Lamport, LaTeX User’s Guide and Reference Manual, 2nd edition, Addison-Wesley, Reading, Massachusetts, 1994.

[2] Michel Goossens, Frank Mittelbach, and Alexander Samarin, The L^{A}TEX Companion, Addison-Wesley, Reading, Massachusetts, 1994.

[3] Donald Ervin Knuth, The TEXbook, Addison-Wesley, Reading, Massachusetts, 1986.

[4] American Mathematical Society and Michael Downes, Technical notes on the amsmath package Version 2.0, 1999/10/29. (Available from CTAN as file

technote.tex.)

[5] Frank Mittelbach, Rainer Schöpf, Michael Downes, and David M. Jones, The amsmath package Version 2.13, 2000/07/18. (Available from CTAN as file

amsmath.dtx.)

[6] Timothy Van Zandt, The fancybox package Version 1.3, 2000/09/19 (Available from CTAN as file fancybox.sty.)

[7] Sebastian Rahtz, Hypertext marks in L^{A}TEX Version v6.74m, 2003/11/30.

(Available from http://www.tug.org/applications/hyperref as file
hyperref.dtx.)

[8] David Carlisle, The keyval Package, Version 1.13, 1999/03/16. (Available from CTAN as file keyval.dtx.)

[9] Kresten Krab Thorup, Frank Jensen, and Chris Rowley. The calc package, Version 4.1b, 1998/07/07, (Available from CTAN as file calc.dtx.)

[10] Wolfgang May and Andreas Schlechte, The ntheorem package Version 1.203, 2002/01/07, (Available from CTAN as file ntheorem.dtx.)

[11] David Carlisle. The showkeys package, v3.12, 1997/06/12, (Available from

    CTAN as file showkeys.dtx.)